OpenStudy (anonymous):

Solve each system. In your work, identify the shape of the graph of each equation. 7x^2 + x^2 = 64 x + y = 4

OpenStudy (anonymous):

Thanks! Can you help me?

OpenStudy (misty1212):

\[7x^2 + x^2 = 64\\ x + y = 4\]

OpenStudy (misty1212):

i am going to assume it is a typo and really was \[7x^2 + y^2 = 64 \\x + y = 4\]

OpenStudy (misty1212):

if not please say so otherwise \[x+y=4\iff y=4-x\] solve \[7x^2+(4-x)^2=6\]

OpenStudy (misty1212):

@DanJS not to be argumentative, but \[ax^2+by^2=c\] is certainly not a circle, less off course \(a=b\)

OpenStudy (misty1212):

i would go ahead and solve \[7x^2+(4-x)^2=6\] not worrying too much about the shapes

OpenStudy (misty1212):

then again it does not have a solution can you check the question again?

OpenStudy (anonymous):

@misty1212 the equations you posted are correct.

OpenStudy (misty1212):

\[7x^2+(4-x)^2=64\]

OpenStudy (misty1212):

solve that one, it is not hard, do you know how to do it?

OpenStudy (anonymous):

not really

OpenStudy (misty1212):

multiply out first on the left

OpenStudy (misty1212):

\[7x^2+x^2-8x+16=64\] or \[8x^2-8x-48=0\]

OpenStudy (misty1212):

divide all by 8 and get to \[x^2-x-6=0\] then factor

OpenStudy (anonymous):

would it be \[(x - 3)(x + 2)?\]

OpenStudy (misty1212):

yes

OpenStudy (misty1212):

so \(x=3,x=-2\) then find the corresponding \(y\) values

OpenStudy (anonymous):

How'd you "multiply out"? Where'd the \[x^2, 8x, and 16 come from?\]

OpenStudy (misty1212):

\[(4-x)^2=(4-x)(4-x)\] etc

OpenStudy (anonymous):

Ohhhhh okay. How do i find the "corresponding y- values?"

OpenStudy (misty1212):

plug in the x values in to the part that says \(x+y=4\) and see what you get for \(y\)

OpenStudy (anonymous):

\[(3, 1) (-2, 6)\]